1) the two sequences of numbers {1,4,16,64,.............} and


{3,12,48,192,....................} are mixed as {1,3,4,12,16,48,64,192,.......}.If one number in the mixed series is 1048576 then its just preceeding number is.


(A)262144  (B)786432  (C)3245728  (D) NOT POSSIBLE TO FIND IT


 


 


 


 

Grade:12

1 Answers

Viraj Jorapur
14 years ago
It can be seen that the first sequence is 4^n-1 where n=1,2,3,.... Second sequence is 3*4^n-1 where n=1,2,3,..... 1048576=4^10. So this is the 11th term of the first series. The preceeding term will be the 10th term of the second series. So substitute n=10 in the second series, we get 3*4^9 which will evidently be option B.